OpenStudy (anonymous):

can someone show me how to do this? o.O f(x)=x^2+4 f(x+1)-f(1)

OpenStudy (anonymous):

sub the expression x+1 into the variable in the function, and then subtract from that the function with 1 subbed in for the variable

OpenStudy (anonymous):

F(x+1)-F(x)=(x+1)^2+4 - [1^2+4] and then simplify

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

=x^2+2x+1+4-5=x^2+2x

OpenStudy (anonymous):

how do i do it with a fraction? \[f(x-5)-f(x)/5\]

OpenStudy (anonymous):

function is defined by the same rule?

OpenStudy (anonymous):

just divide the function by 5, e.g., \[(x-5)^2+4-\frac{x^2+4}{5}=...\]now simplify

OpenStudy (anonymous):

\[=x^2-10x+25+4-\frac{1}{5}x^2+\frac{4}{5}=\frac{4}{5}x^2-\frac{46}{5}x+4\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wait, the constant of 4 should be 29 (as in 25+4); sorry someone was talking to me as I was finishing your solution.

OpenStudy (anonymous):

no wait-- its worse than that give a a minute to fix this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[=\frac{4}{5}x^2-10x+29+\frac{4}{5}=\frac{4}{5}x^2-10x+29+\frac{145}{5}+\frac{4}{5}\]\[=\frac{4}{5}x^2-10x+\frac{149}{5}\]this is the correct one (I hope!)

OpenStudy (anonymous):

is that simplified?

OpenStudy (anonymous):

yes there are no more like terms, and i think i got it right this time